# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
and treat $RGe_{z}$ as a superalgebra in the appropriate way. (Note that unlike in 3.10 we are yet not assuming that $G_{\bar{0}}$ is a proper subgroup of $G$.)
As stated in [22, (4.3)], we have the following isomorphisms of $R$-superalgebras (see 2.7 for the notation $\mathcal{T}_{n}^{R}$):
$\begin{split}R\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$ (5.2)
Remark 5.2.
If $p\equiv 1~{\rm mod}~4$, then $\mathbb{Z}_{p}$ (resp. $\mathbb{F}_{p}$) contains a primitive $4$-th root of unity. Since $R$ is a flat $R_{p}$-module (see [19, Lemma 4.1]), the $R$-superalgebra isomorphisms (5.2) restrict to isomorphisms of $R_{p}$-superalgebras
$\begin{split}R_{p}\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n}^{R_{p}},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R_{p}\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n}^{R_{p}},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$
If $p\equiv 3~{\rm mod}~4$, we still have the first $R_{p}$-superalgebra isomorphism but the second does not hold, because this time $R_{p}$ does not contain any primitive $4$-th root of unity.
5.3.
@@
As in [22], if $X\subseteq[n]$, then $S_{X}$ will signify the subgroup of $S_{n}$ consisting of all elements that fix $\X[n]\backslash$ X pointwise and $\widetilde{S}_{X}$ will denote $\pi_{n}^{-1}(S_{X})\leq\widetilde{S}_{n}$. Moreover, for disjoint $X_{1},\cdots,X_{i}\subseteq[n]$, let $\widetilde{S}_{X_{1},\cdots,X_{r}}$ be the subgroup of $\widetilde{S}_{n}$ generated by $\widetilde{S}_{X_{1}},\cdots,\widetilde{S}_{X_{i}}$. In particular,
$\widetilde{S}_{X_{1},\cdots,X_{i}}\cong\widetilde{S}_{X_{1}}\times_{z}\cdots\times_{z}\widetilde{S}_{X_{i}}$
@@
$R\widetilde{S}_{X_{1},\cdots,X_{i}}e_{z}\cong R\widetilde{S}_{X_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{X_{i}}e_{z}\cong\mathcal{T}_{|X_{1}|}^{R}\otimes_{R}\cdots\otimes_{R}\mathcal{T}_{|X_{i}|}^{R}$ (5.3)
@@
$\widetilde{A}_{X}:=\widetilde{S}_{X}\cap\widetilde{A}_{n}\quad\text{and}\quad\widetilde{A}_{X_{1},\cdots,X_{i}}:=\widetilde{S}_{X_{1},\cdots,X_{i}}\cap\widetilde{A}_{n}.$
@@
$\widetilde{S}_{n_{1},\cdots,n_{i}}:=\widetilde{S}_{Y_{1},\cdots,Y_{i}}\cong\widetilde{S}_{n_{1}}\times_{z}\cdots\times_{z}\widetilde{S}_{n_{i}}\leq\widetilde{S}_{n}\quad\text{and}\quad\widetilde{A}_{n_{1},\cdots,n_{i}}=\widetilde{S}_{n_{1},\cdots,n_{i}}\cap\widetilde{A}_{n}\leq\widetilde{A}_{n}.$
If some of the $n_{l}$’s are equal to each other, we use the exponential notation, as for example for the standard Young subgroup $\widetilde{S}_{r+ip,p^{d-i}}=\widetilde{S}_{r+ip,p,\cdots,p}$.
Following Schur [32], to every strict $\lambda\in\mathscr{P}_{0}(n)$, one associated canonically a spin character $\chi_{\lambda}$ of $\widetilde{S}_{n}$ such that $\chi_{\lambda}$ is irreducible if $n-h(\lambda)$ is even and $\chi_{\lambda}=\chi_{\lambda}^{+}+\chi_{\lambda}^{-}$ for irreducible characters $\chi_{\lambda}^{\pm}$ if $n-h(\lambda)$ is odd. Moreover,
@@
is a complete irredundant set of irreducible spin characters of $\widetilde{S}_{n}$.
Theorem 5.4 ([17], see also [21, Theorem 5.2.4]).
@@
The blocks of $R\widetilde{S}_{n}$ is described by the following theorem:
Theorem 5.5 ([17, Theorem 1.1], see also [22, Theorem 4.3]).
@@
    The spin blocks of $R\widetilde{S}_{n}$ are labelled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core and $d$ is a non-negative integer with $n=|\rho|+dp$. If $d=0$ and $n-h(\rho)$ is odd then $(\rho,0)$ labels two spin blocks of $R\widetilde{S}_{n}$ with corresponding block idempotents $e_{\rho,0}^{+},e_{\rho,0}^{-}\in R\widetilde{S}_{n}e_{z}$. In all cases we denote the block idempotent (or sum of two block idempotents) corresponding to $(\rho,d)$ by $e_{\rho,d}\in R\widetilde{S}_{n}e_{z}$.
@@
    The character $\chi_{\lambda}^{\pm}$ lies in the (sum of) block(s) $\mathcal{O}\widetilde{S}_{n}e_{\rho,d}$, where $\rho$ is the $\bar{p}$-core of $\lambda$. If $d=0$ and $\rho$ is odd, we choose the labeling so that $\chi_{\rho}^{\pm}$ lies in the block $\mathcal{O}\widetilde{S}_{n}e_{\rho,0}^{\pm}$.
Defect groups and Brauer correspondents of blocks of $R\widetilde{S}_{n}$ are described as follows:
Theorem 5.6 ([5, Theorem A and Corollary 26], see also [22, Theorem 4.3]).
@@
    If $d=0$, then each block of $R\widetilde{S}_{n}e_{\rho,d}$ has trivial defect group. If $d>0$, then any Sylow $p$-subgroup $D$ of $\widetilde{S}_{V}\leq\widetilde{S}_{n}$ is a defect group of $R\widetilde{S}_{n}e_{\rho,d}$.
@@
    If $d=0$, then each block of $R\widetilde{S}_{n}e_{\rho,d}$ is its own Brauer correspondent. If $d>0$, then, setting $D$ to be as in (i), the Brauer correspondent of $R\widetilde{S}_{n}e_{\rho,d}$ in $N_{\widetilde{S}_{n}}(D)$ is $RN_{\widetilde{S}_{n}}(D)e_{\rho,0}$. Here we view $e_{\rho,0}$ as an element of $RN_{\widetilde{S}_{n}}(D)$ via $e_{\rho,0}\in R\widetilde{S}_{U}\hookrightarrow RN_{\widetilde{S}_{n}}(D)$. If $\rho=\varnothing$, we interpret $e_{\varnothing,0}$ as $e_{z}$.
Remark 5.7 ([22, Remark 4.4]).
The $R\widetilde{S}_{n}e_{\rho,d}$’s are precisely the superblocks of $R\widetilde{S}_{n}e_{z}$. That is, the finest decomposition of $R\widetilde{S}_{n}e_{z}$ into a direct sum of two-sided ideals that are invariant under the involution $\sigma_{R\widetilde{S}_{n}}$ (see (2.1) for this notation). This follows from the distribution of characters given in Theorem 5.5 (ii). In particular, we always have $e_{\rho,d}\in R\widetilde{A}_{n}e_{z}$.
The spin blocks of $R\widetilde{A}_{n}$ are described in [18, Proposition 3.16]:
@@
For each $n>1$, $\bar{p}$-core $\rho$ and $d$ a non-negative integer with $n=|\rho|+dp$, $R\widetilde{A}_{n}e_{\rho,d}$ is a single block of $R\widetilde{A}_{n}$ unless $d=0$ and $\rho$ is even. In this latter case $R\widetilde{A}_{n}e_{\rho,d}$ is a direct sum of two blocks of $R\widetilde{A}_{n}$. If $n=1$, then $R\widetilde{A}_{1}e_{(1),0}=R\widetilde{A}_{1}e_{(1),0}=Re_{z}$ is a single block of $R\widetilde{A}_{1}$.
If $d=0$, once again, the defect group is trivial. In all other cases the defect group of $R\widetilde{A}_{n}e_{\rho,d}$ is the same as that of $R\widetilde{S}_{n}e_{\rho,d}$.
@@
$B^{\rho,d}_{R}:=R\widetilde{S}_{n}e_{\rho,d}~~{\rm and}~~(B^{\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}.$
It is easy to see that $R\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R\widetilde{S}_{n}e_{\rho,d}$. We will often assume $d>0$ to ensure that blocks and superblocks coincide. However, the corresponding statements for trivial defect blocks are completely elementary.
@@
Keep the notation of Theorem 5.5. The central idempotent $e_{\rho,d}$ of $R\widetilde{S}_{n}$ is contained in the subring $R_{p}\widetilde{S}_{n}$, hence also contained in $R_{p}\widetilde{A}_{n}$ by Remark 5.7.
This proposition can be deduced from the following lemma, which is a slight generalisation of [16, Proposition 2.2].
@@
$B^{\rho,d}_{R_{p}}:=R_{p}\widetilde{S}_{n}e_{\rho,d}~~{\rm and}~~(B^{\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}\widetilde{A}_{n}e_{\rho,d}.$
As before, it is easy to see that $R_{p}\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R_{p}\widetilde{S}_{n}e_{\rho,d}$.
@@
Let $D_{i}$ is a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$. With this notation, the following now follows easily from Theorem 5.5:
Lemma 5.11 ([22, Lemma 4.6]).
Let $D$ be a defect group of the block algebra $R\widetilde{S}_{n}e_{\rho,d}$. Then $D$ is abelian if and only if $d<p$. In this case we can choose $D=D_{1}\times\dots\times D_{d}$.
5.12.
Keep the notation in (5.4). Let $\iota:S_{d}\to S_{V}$ be the natural permutation action of $S_{d}$ on the set $\{P_{1},\cdots,P_{d}\}$: more precisely, for any $1\leq i\leq d$, $1\leq j\leq p$ and $w\in S_{d}$
@@
For each $w\in S_{d}$, choose a lift $T_{w}$ of $\iota(w)$ to $\widetilde{S}_{V}$, and set $T_{i}:=T_{s_{i}}$ for all $1\leq i\leq d-1$.
Note that every Sylow $p$-subgroup of $S_{V_{i}}$ lifts uniquely and isomorphically to a Sylow $p$-subgroup of $\widetilde{S}_{V_{i}}$. Hence we can choose the $D_{i}$’s such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $1\leq i\leq d$ and $w\in S_{d}$. Indeed, one can first construct Sylow $p$-subgroups of each of the $S_{V_{i}}$’s that get permuted by the $\iota(w)$’s and then lift to $\widetilde{S}_{V_{i}}$. By the construction of the $T_{w}$’s we also have $T_{w}\widetilde{S}_{V_{i}}T_{w}^{-1}=\widetilde{S}_{V_{w(i)}}$, for all $1\leq i\leq d$ and $w\in S_{d}$.
Note that $R\widetilde{S}_{V_{1}}e_{z}\cong R\widetilde{S}_{p}e_{z}$ via $t_{r+m}e_{z}\mapsto t_{m}e_{z}$. We now fix isomorphisms between the $R\widetilde{S}_{V_{i}}e_{z}$’s. For each $1<i\leq d$, we identify $R\widetilde{S}_{V_{i}}e_{z}$ with $R\widetilde{S}_{V_{1}}e_{z}\cong R\widetilde{S}_{p}e_{z}$ via the isomorphism
$R\widetilde{S}_{V_{1}}e_{z}\to R\widetilde{S}_{V_{i}}e_{z},\quad a\mapsto(-1)^{i|a|}T_{(1,i)}aT_{(1,i)}^{-1}.$ (5.5)
(This does not depend on our choice of $T_{(1,i)}$, because $T_{(1,i)}$ is uniquely determined up to multiplication by $z$.) Through these isomorphisms and using (5.3), we can identify the superalgebra
$R\widetilde{S}_{V_{1},\cdots,V_{d}}e_{z}\cong R\widetilde{S}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}_{V_{d}}e_{z}$ (5.6)
with $(R\widetilde{S}_{p}e_{z})^{\otimes d}$. Clearly (5.5) and (5.6) still hold with $R$ replaced by $R_{p}$ - because all coefficients involved in these isomorphisms are contained in $R_{p}$.
5.13.

### v5

@@ 6 On spin blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$
and treat $RGe_{z}$ as a superalgebra in the appropriate way. (Note that unlike in 4.10 we are yet not assuming that $G_{\bar{0}}$ is a proper subgroup of $G$.)
As stated in [23, (4.3)], we have the following isomorphisms of $R$-superalgebras (see 2.7 for the notation $\mathcal{T}_{n,R}$):
$\begin{split}R\widetilde{S}_{n}^{+}e_{z}&\cong\mathcal{T}_{n,R},~t_{i}e_{z}\mapsto\tau_{i},\ ze_{z}\mapsto-1\\ R\widetilde{S}_{n}^{-}e_{z}&\cong\mathcal{T}_{n,R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1.\end{split}$ (6.2)
6.2.
@@
As in [23], if $X\subseteq[n]$, then $S_{X}$ will signify the subgroup of $S_{n}$ consisting of all elements that fix $\X[n]\backslash$ X pointwise and $\widetilde{S}^{\eta}_{X}$ will denote $\pi_{n}^{-1}(S_{X})\leq\widetilde{S}^{\eta}_{n}$. Moreover, for disjoint $X_{1},\cdots,X_{i}\subseteq[n]$, let $\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}$ be the subgroup of $\widetilde{S}^{\eta}_{n}$ generated by $\widetilde{S}^{\eta}_{X_{1}},\cdots,\widetilde{S}^{\eta}_{X_{i}}$. In particular,
$\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}\cong\widetilde{S}^{\eta}_{X_{1}}\times_{z}\cdots\times_{z}\widetilde{S}^{\eta}_{X_{i}}$
@@
$R\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}e_{z}\cong R\widetilde{S}^{\eta}_{X_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}^{\eta}_{X_{i}}e_{z}$ (6.3)
@@
$\widetilde{A}_{X}:=\widetilde{S}^{\eta}_{X}\cap\widetilde{A}_{n}\quad\text{and}\quad\widetilde{A}_{X_{1},\cdots,X_{i}}:=\widetilde{S}^{\eta}_{X_{1},\cdots,X_{i}}\cap\widetilde{A}_{n}.$
@@
$\widetilde{S}^{\eta}_{n_{1},\cdots,n_{i}}:=\widetilde{S}^{\eta}_{Y_{1},\cdots,Y_{i}}\cong\widetilde{S}^{\eta}_{n_{1}}\times_{z}\cdots\times_{z}\widetilde{S}^{\eta}_{n_{i}}\leq\widetilde{S}^{\eta}_{n}\quad\text{and}\quad\widetilde{A}_{n_{1},\cdots,n_{i}}=\widetilde{S}^{\eta}_{n_{1},\cdots,n_{i}}\cap\widetilde{A}_{n}\leq\widetilde{A}_{n}.$
If some of the $n_{l}$’s are equal to each other, we use the exponential notation, as for example for the standard Young subgroup $\widetilde{S}^{\eta}_{r+ip,p^{d-i}}=\widetilde{S}^{\eta}_{r+ip,p,\cdots,p}$.
Following Schur [32], to every strict $\lambda\in\mathscr{P}_{0}(n)$, one associated canonically a spin character $\chi_{\lambda}$ of $\widetilde{S}^{\eta}_{n}$ such that $\chi_{\lambda}$ is irreducible if $n-h(\lambda)$ is even and $\chi_{\lambda}=\chi_{\lambda}^{+}+\chi_{\lambda}^{-}$ for irreducible characters $\chi_{\lambda}^{\pm}$ if $n-h(\lambda)$ is odd. Moreover,
@@
is a complete irredundant set of irreducible spin characters of $\widetilde{S}^{\eta}_{n}$.
Theorem 6.3 ([18], see also [22, Theorem 5.2.4]).
@@
The blocks of $R\widetilde{S}^{\eta}_{n}$ is described by the following theorem:
Theorem 6.4 ([18, Theorem 1.1], see also [23, Theorem 4.3]).
@@
    The spin blocks of $R\widetilde{S}^{\eta}_{n}$ are labelled by pairs $(\rho,d)$, where $\rho$ is a $\bar{p}$-core and $d$ is a non-negative integer with $n=|\rho|+dp$. If $d=0$ and $n-h(\rho)$ is odd then $(\rho,0)$ labels two spin blocks of $R\widetilde{S}^{\eta}_{n}$ with corresponding block idempotents $e_{\rho,0}^{+},e_{\rho,0}^{-}\in R\widetilde{S}^{\eta}_{n}e_{z}$. In all cases we denote the block idempotent (or sum of two block idempotents) corresponding to $(\rho,d)$ by $e_{\rho,d}\in R\widetilde{S}^{\eta}_{n}e_{z}$.
@@
    The character $\chi_{\lambda}^{\pm}$ lies in the (sum of) block(s) $\mathcal{O}\widetilde{S}^{\eta}_{n}e_{\rho,d}$, where $\rho$ is the $\bar{p}$-core of $\lambda$. If $d=0$ and $\rho$ is odd, we choose the labeling so that $\chi_{\rho}^{\pm}$ lies in the block $\mathcal{O}\widetilde{S}^{\eta}_{n}e_{\rho,0}^{\pm}$.
Defect groups and Brauer correspondents of blocks of $R\widetilde{S}^{\eta}_{n}$ are described as follows:
Theorem 6.5 ([6, Theorem A and Corollary 26], see also [23, Theorem 4.3]).
@@
    If $d=0$, then each block of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$ has trivial defect group. If $d>0$, then any Sylow $p$-subgroup $D$ of $\widetilde{S}^{\eta}_{V}\leq\widetilde{S}^{\eta}_{n}$ is a defect group of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
@@
    If $d=0$, then each block of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$ is its own Brauer correspondent. If $d>0$, then, setting $D$ to be as in (i), the Brauer correspondent of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$ in $N_{\widetilde{S}^{\eta}_{n}}(D)$ is $RN_{\widetilde{S}^{\eta}_{n}}(D)e_{\rho,0}$. Here we view $e_{\rho,0}$ as an element of $RN_{\widetilde{S}^{\eta}_{n}}(D)$ via $e_{\rho,0}\in R\widetilde{S}^{\eta}_{U}\hookrightarrow RN_{\widetilde{S}^{\eta}_{n}}(D)$. If $\rho=\varnothing$, we interpret $e_{\varnothing,0}$ as $e_{z}$.
Remark 6.6 ([23, Remark 4.4]).
The $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$’s are precisely the superblocks of $R\widetilde{S}^{\eta}_{n}e_{z}$. That is, the finest decomposition of $R\widetilde{S}^{\eta}_{n}e_{z}$ into a direct sum of two-sided ideals that are invariant under the involution $\sigma_{R\widetilde{S}^{\eta}_{n}}$ (see (2.1) for this notation). This follows from the distribution of characters given in Theorem 6.4 (ii). In particular, we always have $e_{\rho,d}\in R\widetilde{A}_{n}e_{z}$.
The spin blocks of $R\widetilde{A}_{n}$ are described in [19, Proposition 3.16]:
@@
For each $n>1$, a $\bar{p}$-core $\rho$ and a non-negative integer $d$ with $n=|\rho|+dp$, $R\widetilde{A}_{n}e_{\rho,d}$ is a single block of $R\widetilde{A}_{n}$ unless $d=0$ and $\rho$ is even. In this latter case $R\widetilde{A}_{n}e_{\rho,d}$ is a direct sum of two blocks of $R\widetilde{A}_{n}$. If $n=1$, then $R\widetilde{A}_{1}e_{(1),0}=R\widetilde{A}_{1}e_{(1),0}=Re_{z}$ is a single block of $R\widetilde{A}_{1}$.
If $d=0$, once again, the defect group is trivial. In all other cases the defect group of $R\widetilde{A}_{n}e_{\rho,d}$ is the same as that of $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
@@
$B^{\eta,\rho,d}_{R}:=R\widetilde{S}^{\eta}_{n}e_{\rho,d}~~{\rm and}~~(B^{\eta,\rho,d}_{R})_{\bar{0}}:=R\widetilde{A}_{n}e_{\rho,d}.$
It is easy to see that $R\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$. We will often assume $d>0$ to ensure that blocks and superblocks coincide. However, the corresponding statements for trivial defect blocks are completely elementary.
@@
Keep the notation of Theorem 6.4. The central idempotent $e_{\rho,d}$ of $R\widetilde{S}^{\eta}_{n}$ is contained in the subring $R_{p}\widetilde{S}^{\eta}_{n}$, hence also contained in $R_{p}\widetilde{A}_{n}$ by Remark $6.$
This proposition can be deduced from the following lemma, which is a slight generalisation of [17, Proposition 2.2].
@@
$B^{\eta,\rho,d}_{R_{p}}:=R_{p}\widetilde{S}^{\eta}_{n}e_{\rho,d}~~{\rm and}~~(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}:=R_{p}\widetilde{A}_{n}e_{\rho,d}.$
As before, it is easy to see that $R_{p}\widetilde{A}_{n}e_{\rho,d}$ is indeed the purely even subalgebra of the superalgebra $R_{p}\widetilde{S}^{\eta}_{n}e_{\rho,d}$.
@@
Let $D_{i}$ is a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$. With this notation, the following lemma follows easily from Theorem 6.4:
Lemma 6.10 ([23, Lemma 4.6]).
Let $D$ be a defect group of the block algebra $R\widetilde{S}^{\eta}_{n}e_{\rho,d}$. Then $D$ is abelian if and only if $d<p$. In this case we can choose $D=D_{1}\times\dots\times D_{d}$.
6.11.
Keep the notation in (6.4). Let $\iota:S_{d}\to S_{V}$ be the natural permutation action of $S_{d}$ on the set $\{V_{1},\cdots,V_{d}\}$: more precisely, for any $1\leq i\leq d$, $1\leq j\leq p$ and $w\in S_{d}$
@@
For each $w\in S_{d}$, choose a lift $T_{w}$ of $\iota(w)$ to $\widetilde{S}^{\eta}_{V}$, and set $T_{i}:=T_{s_{i}}$ for all $1\leq i\leq d-1$.
Note that every Sylow $p$-subgroup of $S_{V_{i}}$ lifts uniquely and isomorphically to a Sylow $p$-subgroup of $\widetilde{S}^{\eta}_{V_{i}}$. Hence we can choose the $D_{i}$’s such that $T_{w}D_{i}T_{w}^{-1}=D_{w(i)}$, for all $1\leq i\leq d$ and $w\in S_{d}$. Indeed, one can first construct Sylow $p$-subgroups of each of the $S_{V_{i}}$’s that get permuted by the $\iota(w)$’s and then lift to $\widetilde{S}^{\eta}_{V_{i}}$. By the construction of the $T_{w}$’s we also have $T_{w}\widetilde{S}^{\eta}_{V_{i}}T_{w}^{-1}=\widetilde{S}^{\eta}_{V_{w(i)}}$, for all $1\leq i\leq d$ and $w\in S_{d}$.
Note that $R\widetilde{S}^{\eta}_{V_{1}}e_{z}\cong R\widetilde{S}^{\eta}_{p}e_{z}$ via $t_{r+m}e_{z}\mapsto t_{m}e_{z}$. We now fix isomorphisms between the $R\widetilde{S}^{\eta}_{V_{i}}e_{z}$’s. For each $1<i\leq d$, we identify $R\widetilde{S}^{\eta}_{V_{i}}e_{z}$ with $R\widetilde{S}^{\eta}_{V_{1}}e_{z}\cong R\widetilde{S}^{\eta}_{p}e_{z}$ via the isomorphism
$R\widetilde{S}^{\eta}_{V_{1}}e_{z}\to R\widetilde{S}^{\eta}_{V_{i}}e_{z},\quad a\mapsto(-1)^{i|a|}T_{(1,i)}aT_{(1,i)}^{-1}.$ (6.5)
(This does not depend on the choice of $T_{(1,i)}$, because $T_{(1,i)}$ is uniquely determined up to multiplication by $z$.) Through these isomorphisms and using (6.3), we can identify the superalgebra
$R\widetilde{S}^{\eta}_{V_{1},\cdots,V_{d}}e_{z}\cong R\widetilde{S}^{\eta}_{V_{1}}e_{z}\otimes_{R}\cdots\otimes_{R}R\widetilde{S}^{\eta}_{V_{d}}e_{z}$ (6.6)
with $(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$. Clearly (6.3), (6.5) and (6.6) still hold with $R$ replaced by $R_{p}$ - because all coefficients involved in these isomorphisms are contained in $R_{p}$.
Lemma 6.12 (a slight refinement of [23, Lemma 4.7]).
Keep the notation in 6.11. Up to modifying the choice of $T_{i}$ for $1\leq i\leq d-1$, the set $\{T_{i}\mid 1\leq i\leq d-1\}$ satisfies the following equalities:
  1. 1.
    $),(T_{i}e_{z})(a_{1}\otimes\cdots\otimes a_{d})(T_{i}^{-1}e_{z})=(-1)^{\sum_{j\neq i,i+1}|a_{j}|}({}^{s_{i}}(a_{1}\otimes\cdots\otimes a_{d}$)), for all $1\leq i\leq d-1$ and $a_{1}\otimes\dots\otimes a_{d}\in(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$.
  2. 2.
    $(T_{i}e_{z})^{2}=\eta(-1)^{\frac{(p-1)}{2}}e_{z},$ for all $1\leq i\leq d-1$.
  3. 3.
    $(T_{i}e_{z})(T_{l}e_{z})=-(T_{l}e_{z})(T_{i}e_{z}),$ for all $1\leq i,l\leq d-1$ with |il|>1|i-l|>1.
  4. 4.
    $(T_{i}e_{z}T_{i+1}e_{z})^{3}=(T_{i+1}e_{z}T_{i}e_{z})^{3},$ for all $1\leq i\leq d-2$.
The result also holds for $R_{p}$ instead of $R$.
Proof. Equality (i) does not depend on the choice of $T_{i}$’s. Indeed, if $a_{1}\otimes\cdots\otimes a_{d}\in(R\widetilde{S}^{\eta}_{p}e_{z})^{\otimes d}$, we have
$(zT_{i}e_{z})(a_{1}\otimes\cdots\otimes a_{d})(z^{-1}T_{i}^{-1}e_{z})=T_{i}(a_{1}\otimes\cdots\otimes a_{d})T_{i}^{-1}$
$=$ $T_{(1,i)}T_{(1,i+1)}T_{(1,i)}(a_{1}\otimes\cdots\otimes a_{d})T_{(1,i)}^{-1}T_{(1,i+1)}^{-1}T_{(1,i)}^{-1}.$
Equality (i) now follows from a direct calculation using (6.5).
Equality (iii) also does not depend on the $T_{i}$’s. Noting that $ze_{z}=-e_{z}$, equality (iii) follows immediately from (6.3).
Recall that for each $1\leq i\leq d-1$, $T_{i}$ is a lift of $\iota(s_{i})$, where
$\iota(s_{i})=(r+p(i-1)+1,r+pi+1)(r+p(i-1)+2,r+pi+2)\cdots(r+pi,r+pi+p).$
So $T_{i}^{2}=z^{p(p-1)/2}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{+}$, and $T_{i}^{2}=z\cdot z^{p(p-1)/2}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{-}$. Noting that $ze_{z}=-e_{z}$, we have $(T_{i}e_{z})^{2}=(-1)^{p(p-1)/2}e_{z}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{+}$, and $(T_{i}e_{z})^{2}=(-1)\cdot(-1)^{p(p-1)/2}e_{z}$ if $\widetilde{S}^{\eta}_{n}=\widetilde{S}_{n}^{-}$. In other words, we have $(T_{i}e_{z})^{2}=\eta(-1)^{\frac{(p-1)}{2}}e_{z}$. This proves equality (ii). Again, equality (ii) does not depend on the choice of $T_{i}$.
Finally, since the $\pi_{n}(T_{i})$’s satisfy the braid relations, property (iv) holds up to a $z$, for each $1\leq i\leq d-2$. We may therefore change the choice of some of $T_{i}$’s to ensure that equality (iv) holds.
[…]
6.13.